OpenStudy (anonymous):

the sum of two mumbers is 28 and the difference is 4. what are the numbers????

OpenStudy (anonymous):

solve the system: x+y = 28 x-y = 4

hero (hero):

2nd Hint: Add both equations together to eliminate one of the variables

OpenStudy (anonymous):

sum is 28 so x+y= 28

OpenStudy (anonymous):

let x be the bigger number let y be the smaller number

OpenStudy (anonymous):

x+y=28